Bibliography

box

23/03-04 16:20

[berndt1995a] Hendrik Berndt and Peter Graubmann. Service architecture, service session and service federation in TINA-C. In International Switching Symposium, ISS’95, Berlin, Germany, April 1995.

[brown1995b] Dave Brown and Stefano Montesi. Requirements upon TINA-C architecture. TINA Baseline TB-MH.002-2.0-94, TINA, February 1995.

[chapman1995a] Martin Chapman and Stefano Montesi. Overall concepts and principles of TINA. TINA Baseline TB-MDC.018-1.0-94, TINA, February 1995.

[desoto1997a] Alden DeSoto. Using the Beans development kit 1.0. Technical report, JavaSoft, September 1997.

[graubman1994a] P. Graubman, W. Hwang, M. Kudela, K. MacKinnon, and N. Mercouroff. Engineering modelling concepts (DPE architecture). TINA Baseline TB-NS.005-2.0-94, TINA, December 1994.

[hamilton1997a] Graham Hamilton. JavaBeans. Api specification, Sun Microsystems, July 1997. Version 1.01.

[iso-iec1995a] ISO/IEC. Open distributed processing reference model, part 1: Overview. ITU-T Rec. X.901 — ISO/IEC 10746-1, ISO/IEC, 1995.

[iso-iec1995b] ISO/IEC. Open distributed processing reference model, part 2: Foundations. ITU-T Rec. X.902 — ISO/IEC 10746-2, ISO/IEC, 1995.

[iso-iec1995c] ISO/IEC. Open distributed processing reference model, part 3: Architecture. ITU-T Rec. X.903 — ISO/IEC 10746-3, ISO/IEC, 1995.

[iso-iec1995d] ISO/IEC. Open distributed processing reference model, part 4: Architectural semantics amendment. ITU-T Rec. X.904 — ISO/IEC 10746-4, ISO/IEC, 1995.

[iso-iec1995e] ISO/IEC. ODP trading function. ITU-T Draft Rec. X.9tr — ISO/IEC DIS 13235, ISO/IEC, 1995.

[iso-iec1998a] ISO/IEC. Working draft for open distributed processing reference model, quality of service. Approved meeting output, editor’s draft 6.4, ISO/IEC, June 1998.

[iso-iec1998b] ISO/IEC. Open distributed processing, interface references and binding. Itu-t draft recommendation x.300, ISO/IEC, September 1998.

[javasoft1996a] JavaSoft. Java core reflection, API and specification. Technical report, Sun Microsystems, Inc., September 1996.

[javasoft1998a] JavaSoft. Java remote method invocation — distributed computing for Java. White paper, Sun Microsystems, Inc., 1998.

[kudela1994a] M. Kudela and K. MacKinnon. Engineering modelling concepts (DPE kernel specifications). TINA Report TR-KMK.001-1.1-94, TINA, November 1994.

[microsoft1996a] Microsoft. DCOM technical overview. Microsoft Windows NT Server white paper, Microsoft Corporation, 1996.

[omg1993a] Object Managment Group. Object request broker architecture. OMG TC Document 93.7.2, Object Managment Group, July 1993. (version 0.0, draft).

[wakano1994a] Masaki Wakano, Motoharu Kawanishi, and Lars Richter. Information model to support TINA service and management applications. In IEEE Global Telecommunications Conference, GLOBECOM’94, San Francisco, California, USA, November 1994.